OpenStudy (anonymous):

Find the indicated nth partial sum of the arithmetic sequence. 1.6, 4.2, 6.8, 9.4, ..., n = 50

OpenStudy (amistre64):

+2.6 ?

OpenStudy (anonymous):

The nth term = dn + c. I think I use this to solve but when I do none of my answers match up with my choices

OpenStudy (anonymous):

and yes 2.6 is the common difference (d)

OpenStudy (amistre64):

1.6 + 2.6(n) ?? think I am wrong...

OpenStudy (anonymous):

my answer choices are: A. 2090 B.3395 C.3265 D.3266 E.3264

OpenStudy (anonymous):

would it not be 1.6+2.6(n-1) ?? Remembering that you start at n=1

OpenStudy (amistre64):

math: thats what I thought, but doubted myself....

OpenStudy (anonymous):

Well thats what I was doing but as you can see none of my answer choices match that answer

OpenStudy (anonymous):

that's because you are looking for the sum of the terms up o the 50th term

OpenStudy (anonymous):

ohhhh wow silly me so I have to go back and add them all up.... sorry about that

OpenStudy (anonymous):

the idea of the question i assume is for you to find the formula which does the sumation for you

OpenStudy (amistre64):

can you do that with an integral of the equation?

OpenStudy (anonymous):

i think to find n, you could use 50=n/2 (1.6+(n-1)2.6)

OpenStudy (anonymous):

no because that would find the area under an axis of real numbers whereas you are looking to sum natural numbers

OpenStudy (anonymous):

\[s \left[ n \right]=n\2(a \left[ 1 \right]+a \left[ n \right])\]

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

no I don't think that is right either.....

OpenStudy (anonymous):

no its not, keep trying though

OpenStudy (anonymous):

Just kidding I figured it out the answer is c

OpenStudy (anonymous):

1.6+49(2.6)=129 50/2(1.6+129)=3265

OpenStudy (anonymous):

oh i see what you meant now by your notation! My mistake, well done :)

OpenStudy (anonymous):

do any of you know or understand how to figure out a problem like this: Find the sum of the infinite series: \[\sum_{i=1}^{\infty}2(-1/4)^i\]

OpenStudy (anonymous):

Thanks It took me a sec to think through it and figure it out lol maybe if I had originally read the question better lol :)

OpenStudy (anonymous):

if i remember correctly there is a formula for the sum of an infinite series, first take 2 out as it is a constant and then apply the formula, (try wikipedia for it because i cannot be sure of my memory). As -1<r<1 you can use the formula where r=-1/4

OpenStudy (anonymous):

Ok thanks